# Semantic Energy: Detecting LLM Hallucination Beyond Entropy

### v1

@@ 1 Introduction
Large Language Models (LLMs) have been widely deployed in various aspects of production and daily life, demonstrating strong capabilities in different fields (Schlegel et al., 2025; Xiang et al., 2025). However, LLMs are still prone to being influenced by hallucinations and are prone to generate incorrect answers in situations where they lack knowledge, thus misleading users into making errors (Zhou et al., 2024; Farquhar et al., 2024). Recently, uncertainty estimation has been shown to be a reliable indicator for detecting hallucinations, reflecting the tendency of an LLM to generate hallucinations (Xiao & Wang, 2021; Huang et al., 2024). When the uncertainty of an LLM response is high, it often suggests a greater likelihood that the response is a hallucination, prompting further actions like self-reflection (Renze & Guven, 2024; Kirchhof et al., 2025), regenerating of answers (Xu et al., 2025), or intervention by human experts (Liu et al., 2025; Hopkins et al., 2025).
@@
A representative case occurs when the model yields identical responses across multiple sampling instances for a given question, as illustrated in Fig. 1. According to semantic entropy, the resulting value is 0, which is considered as a reliable response. However, even answering incorrectly, LLMs might also provide responses with the same semantics. Among samples with consistently semantic responses across multiple responses, the proportion of incorrect responses (like Question3 in Fig. 1) approaches 50% in some datasets. In such cases, it is necessary to leverage the model’s inherent uncertainty for differentiation: even if the LLM provides multiple responses with the same semantics for two different questions, their corresponding reliability still differs. In scenarios with higher inherent uncertainty in the model, the likelihood of the LLM making mistakes is greater.
A substantial amount of previous research has demonstrated that the logits exhibit stronger inherent capabilities in characterizing uncertainty compared to probabilities and the magnitude of logits can indicate whether the model has undergone adequate training in a given scenario (Liu et al., 2020; Fu et al., 2025; Zhang et al., 2024). For example, in out-of-distribution (OOD) detection, studies have highlighted that the logit values for in-distribution (InD) samples are significantly higher than those for OOD samples (Liu et al., 2020). Recent work named LogToKU (Ma et al., 2025) points out that probabilities lose the intensity information of logits during normalization, thus limiting their ability to represent the inherent uncertainty of LLM. From this insight, we propose a new method to improve the failure cases of Semantic Entropy, termed Semantic Energy. Specifically, for a given prompt, we first perform multiple response samplings, followed by semantic sampling. When calculating the final uncertainty, rather than relying on probability as in Semantic Entropy, we estimate the response uncertainty based on logits, enabling the estimated uncertainty to reflect the model’s inherent uncertainty. Our proposed metric significantly outperforms Semantic Entropy in evaluating the reliability of LLM responses, particularly in scenarios where Semantic Entropy fails. The main contributions are as follows:
@@ 2.2 Semantic Entropy and Response Clustering
where $\bar{p}({\bm{x}}^{(i)})$ indicates a normalized distribution over the sampled responses. Due to surface-level variability in language, semantically similar responses may have different forms. Therefore, semantic entropy cluster the responses into $K$ semantically coherent groups:
@@ 3.1 Limitations of Entropy-Based Uncertainty Estimation
The reason behind this is that the entropy calculated based on probabilities only reflects the relative likelihood of a particular LLM response compared to other possible responses generated by the model, rather than the actual probability of that response to the question in the real world. However, when computing the probability of the next token, the model approximates the partition function as the sum of probabilities over its vocabulary. Therefore, for entropy to accurately represent uncertainty, two assumptions must hold: (1) the model has seen all possible responses (i.e., the training distribution matches the real-world distribution perfectly), and (2) the model has fit the training distribution without bias (i.e., the model’s output distribution matches the training distribution exactly). Clearly, neither of these assumptions holds true. Worse still, the LLM response is based on the joint probability prediction of many tokens, rather than the single word classification of traditional discriminative models, which causes the error of the approximated partition function to accumulate to a degree that can no longer be ignored.
@@ 3.2 Energy-Based Confidence Estimation
To address this limitation, we introduce an energy-based formulation that complements semantic entropy and captures epistemic uncertainty. In of physics, lower energy corresponds to a more stable and less random state. Drawing on thermodynamic analogies, we treat lower-energy states as higher-confidence predictions, following the intuition that physical systems evolve toward minimal-energy configurations.
@@ 3.2.1 Boltzmann Distribution
where $k$ is the Boltzmann constant, $T$ is temperature, and $E_{t}^{(i)}$ is the token energy $x_{t}^{(i)}$, and $Z_{t}$ is the partition function. Specifically, for LLMs, $Z_{t}=\sum_{x\in\mathbb{V}}e^{-E_{t}(x)}$ is the normalization value over the entire vocabulary $\mathbb{V}$ (the difference between $\mathcal{V}$ and $\mathbb{V}$ is that $\mathcal{V}$ is the vocabulary in the predefined tokenizer of a specific LLM, while $\mathbb{V}$ is the space of possible next tokens in the real world, which is infinite and intractable.). For simplicity, we assume $Z_{t}$ is constant across $t$. The probability of a complete sequence and the average sequence-level energy can be represented as:
@@ 3.2.2 Specific Implementation in LLMs
where $n$ are the sampling times for every question and a lower energy.
@@
which means that the lower the energy, the lower uncertainty.
@@ 4.1 Setup
We conduct experiments using the Qwen3‑8B model (Yang et al., 2025), and the ERNIE-21B-A3B model (MOE architecture) (Baidu, 2025). Our primary goal is to highlight the differences between probability-based methods and energy-based approaches. Therefore, we use semantic entropy (Farquhar et al., 2024) as the baseline.
@@
Experiments were performed on standard open-domain QA datasets in both Chinese and English: the Chinese dataset CSQA (He et al., 2024) and the English dataset TriviaQA (Joshi et al., 2017). To assess whether the estimated uncertainty can capture the risk that the model makes errors, we estimate the AUROC between uncertainty scores and correctness (i.e., whether the answer is correct).
@@ 4.2 Main Results
Across both models and datasets, semantic energy-based methods consistently outperform semantic entropy. On CSQA, the Boltzmann energy improves AUROC from $71.6\%$} to $76.1\%$} on Qwen3-8B and from $77.4\%$} to $80.2\%$} on ERNIE-21B-A3B. Similar trends are observed on TriviaQA, where Boltzmann energy yields AUROC gains of more than $5\%$ compared to the semantic entropy. Improvements are also reflected in AUPR and FPR@95, indicating better calibration and reduced false positive rates.
These results highlight the robustness of energy-based uncertainty estimation, particularly in low-diversity scenarios where entropy becomes degenerate (details in Table 2). By incorporating internal model states via energy formulations, these methods offer a richer signal for confidence estimation beyond surface-level output variation.
@@ 4.3.1 Results on questions with single cluster
In Table 2, we present the case where all responses share the same semantics, that is, all responses are clustered into a single group as described in the Sec. 3.1. In this scenario, semantic entropy completely fails, whereas semantic energy is still able to provide a certain level of distinction, resulting in semantic energy achieving an average performance improvement of more than 13% compared in terms of AUROC to semantic entropy in cases where the latter is confident. It is important to note that the value of semantic entropy in such cases is always zero, meaning that its performance reflects the expected behavior when the uncertainty indicator is meaningless, for example, the AUPR corresponds to the number of positive samples (i.e. correct responses).
@@ 4.3.2 Results on think Mode
We conduct experiments on Qwen-8B using the CSQA dataset and explore the case where the think mode is enabled. Specifically, we activate the model’s think mode but discard the content within <think>...</think> during evaluation, considering only the response portion. The final results, shown in Fig. 2, are consistent with the observations in Table 1.
@@ 5 Related Work
Recently, numerous uncertainty estimation methods for Large Language Models (LLMs) have been proposed. These include methods that utilize natural language for uncertainty feedback, including heuristically designed and trained approaches (Tao et al., 2025; Xiong et al., 2023; Lin et al., 2023); methods that estimate uncertainty based on model states, including those leveraging prior knowledge or statistical observations of model states (Kostenok et al., 2023; Li et al., 2025; Liu et al., 2024), or observing changes under perturbations (Zhang et al., 2025b; Gao et al., 2024); and methods that take into account the semantics of the response, including consistency-based uncertainty characterizations (Lyu et al., 2025; Bartsch et al., 2023; Xiao et al., 2025) and approaches that integrate semantics with model states (Kuhn et al., 2024; Grewal et al., 2024).

### v2

@@ 1 Introduction
Large Language Models (LLMs) have been widely deployed in various aspects of production and daily life, demonstrating strong capabilities in different fields (Schlegel et al., 2025; Xiang et al., 2025). However, LLMs are still prone to being influenced by hallucinations and are prone to generate incorrect answers in situations where they lack knowledge, thus misleading users into making errors (Zhou et al., 2024; Farquhar et al., 2024). Recently, uncertainty estimation has been shown to be a reliable indicator for detecting hallucinations, reflecting the tendency of an LLM to generate hallucinations (Xiao & Wang, 2021; Huang et al., 2024). When the uncertainty of an LLM response is high, it often suggests a greater likelihood that the response is a hallucination, prompting further actions such as self-reflection (Renze & Guven, 2024; Kirchhof et al., 2025), regenerating of answers (Xu et al., 2025), or intervention by human experts (Liu et al., 2025; Hopkins et al., 2025).
@@
A representative case occurs when the model produces identical responses in multiple sampling instances for a given question, as illustrated in Fig. 1. According to semantic entropy, the resulting value is 0, which is considered a reliable response. However, even answering incorrectly, LLMs might also provide responses with the same semantics. Among samples with consistently semantic responses across multiple responses, the proportion of incorrect responses (like Question3 in Fig. 1) approaches 50% in some datasets. In such cases, it is necessary to leverage the model’s inherent uncertainty for differentiation: even if the LLM provides multiple responses with the same semantics for two different questions, their corresponding reliability still differs. In scenarios with a higher inherent uncertainty in the model, the likelihood of the LLM making mistakes is greater.
Several previous studies have shown that logits exhibit stronger inherent capabilities to characterize uncertainty compared to probabilities, and the magnitude of logits can indicate whether the model has undergone adequate training in a given scenario (Liu et al., 2020; Fu et al., 2025; Zhang et al., 2024). For example, in out-of-distribution (OOD) detection, studies have highlighted that the logit values for in-distribution (InD) samples are significantly higher than those for OOD samples (Liu et al., 2020). Recent work named LogToKU (Ma et al., 2025) points out that probabilities lose the intensity information of logits during normalization, thus limiting their ability to represent the inherent uncertainty of LLM. From this insight, we propose a new method to improve the failure cases of Semantic Entropy, termed Semantic Energy. Specifically, for a given prompt, we first perform multiple response samplings, followed by semantic sampling. When calculating the final uncertainty, rather than relying on probability as in Semantic Entropy, we estimate the response uncertainty based on logits, enabling the estimated uncertainty to reflect the model’s inherent uncertainty. Our proposed metric significantly outperforms Semantic Entropy in evaluating the reliability of LLM responses, particularly in scenarios where Semantic Entropy fails. The main contributions are as follows:
@@ 2.2 Semantic Entropy and Response Clustering
where $\bar{p}({\bm{x}}^{(i)})$ indicates a normalized distribution over the sampled responses. Due to surface-level variability in language, semantically similar responses may have different forms. Therefore, semantic entropy clusters the responses into $K$ semantically coherent groups:
@@ 3.1 Limitations of Entropy-Based Uncertainty Estimation
The reason behind this is that the entropy calculated based on probabilities only reflects the relative likelihood of a particular LLM response compared to other possible responses generated by the model, rather than the actual probability of that response to the question in the real world. However, when computing the probability of the next token, the model approximates the partition function as the sum of probabilities over its vocabulary. Therefore, for entropy to accurately represent uncertainty, two assumptions must hold: (1) the model has seen all possible responses (that is, the training distribution matches the real-world distribution perfectly), and (2) the model has fit the training distribution without bias (that is, the model output distribution matches the training distribution exactly). Clearly, neither of these assumptions holds. Worse still, the LLM response is based on the joint probability prediction of many tokens, rather than the single word classification of traditional discriminative models, which causes the error of the approximated partition function to accumulate to a degree that can no longer be ignored.
@@ 3.2 Energy-Based Confidence Estimation
To address this limitation, we introduce an energy-based formulation that complements semantic entropy and captures epistemic uncertainty. In thermodynamics of physics, lower energy corresponds to a more stable and less random state. Drawing on thermodynamic analogies, we treat lower-energy states as higher-confidence predictions, following the intuition that physical systems evolve toward minimal-energy configurations.
@@ 3.2.1 Boltzmann Distribution
where $k$ is the Boltzmann constant, $T$ is the temperature, and $E_{t}^{(i)}$ is the token energy $x_{t}^{(i)}$, and $Z_{t}$ is the partition function. Specifically, for LLMs, $Z_{t}=\sum_{x\in\mathbb{V}}e^{-E_{t}(x)}$ is the normalization value across the entire vocabulary $\mathbb{V}$ (the difference between $\mathcal{V}$ and $\mathbb{V}$ is that $\mathcal{V}$ is the vocabulary in the predefined tokenizer of a specific LLM, while $\mathbb{V}$ is the space of possible next tokens in the real world, which is infinite and intractable). For simplicity, we assume that $Z_{t}$ is constant across $t$. The probability of a complete sequence and the average sequence-level energy can be represented as:
@@ 3.2.2 Specific Implementation in LLMs
where $n$ are the sampling times for every question.
@@
which means that the lower the energy, the lower the uncertainty.
@@ 4.1 Setup
We conduct experiments using the Qwen3‑8B model (Yang et al., 2025), and the ERNIE-21B-A3B model (MOE architecture) (Baidu, 2025). Our primary goal is to highlight the differences between probability-based methods and energy-based approaches. Therefore, we use the semantic entropy (Farquhar et al., 2024) as a baseline.
@@
Experiments are performed on standard open-domain QA datasets in both Chinese and English: the Chinese dataset CSQA (He et al., 2024) and the English dataset TriviaQA (Joshi et al., 2017). To assess whether the estimated uncertainty can capture the risk that the model makes errors, we estimate the AUROC between uncertainty scores and correctness (i.e., whether the answer is correct).
@@ 4.2 Main Results
In both models and datasets, semantic energy consistently outperforms semantic entropy. On CSQA, the Boltzmann energy improves AUROC from $71.6\%$} to $76.1\%$} on Qwen3-8B and from $77.4\%$} to $80.2\%$} on ERNIE-21B-A3B. Similar trends are observed on TriviaQA, where Boltzmann energy yields AUROC gains of more than $5\%$ compared to the semantic entropy. Improvements are also reflected in AUPR and FPR@95, indicating better calibration and reduced false positive rates.
These results highlight the robustness of energy-based uncertainty estimation, particularly in low-diversity scenarios where entropy becomes degenerate (details in Table 2). By incorporating internal model states via logits, semantic energy captures a richer signal for uncertainty estimation beyond probability-based entropy.
@@ 4.3.1 Results on questions with single cluster
In Table 2, we present the case where all responses share the same semantics, that is, all responses are clustered into a single group as described in Sec. 3.1. In this scenario, semantic entropy completely fails, whereas semantic energy is still able to provide a certain level of distinction, resulting in semantic energy achieving an average performance improvement of more than 13% compared in terms of AUROC to semantic entropy in cases where the latter is confident. It is important to note that the value of semantic entropy in such cases is always zero, meaning that its performance reflects the expected behavior when the uncertainty indicator is meaningless, for example, the AUPR corresponds to the number of positive samples (i.e. correct responses).
@@ 4.3.2 Results on think Mode
We conduct experiments on Qwen-8B using the CSQA dataset and explore the case where the think mode is enabled. Specifically, we activate the think mode but discard the content within <think>...</think> during evaluation, considering only the response portion. The final results, shown in Fig. 2, are consistent with the observations in Table 1.
@@ 5 Related Work
Recently, numerous uncertainty estimation methods for LLMs have been proposed. These include methods that utilize natural language for uncertainty feedback, including heuristically designed and trained approaches (Tao et al., 2025; Xiong et al., 2023; Lin et al., 2023); methods that estimate uncertainty based on model states, including those leveraging prior knowledge or statistical observations of model states (Kostenok et al., 2023; Li et al., 2025; Liu et al., 2024), or observing changes under perturbations (Zhang et al., 2025b; Gao et al., 2024); and methods that take into account the semantics of the response, including consistency-based uncertainty characterizations (Lyu et al., 2025; Bartsch et al., 2023; Xiao et al., 2025) and approaches that integrate semantics with model states (Kuhn et al., 2024; Grewal et al., 2024).